OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

the parallel line to y=-x+6, must have same gradient so it will look like that y=-x+c, lets find c: line y=-x+c passes through (-4,2) so if u subsitute it to equation of searched line y=-x+c, u will get: 2=-(-4)+c 2=4+c 2-4=c c=-2, so y=-x-2